OpenStudy (anonymous):

Question 4 (Multiple Choice Worth 2 points) If a baseball player hits a baseball from 4 feet off the ground with an initial velocity of 64 feet per second, how long will it take the baseball to hit the ground? Use the equation h = –16t2 + 64t + 4. 2 plus or minus square root of 17 end root over 2 quantity of 2 plus or minus square root of 17 all over 2 2 plus or minus 4 square root of 17 quantity of 16 plus or minus square root of 17 all over 2

OpenStudy (anonymous):

Here's what I have: h=-16t^2+64+4 0=-16t^2+64t+4 0=-4(4t^2-16t-1) x=-b+-sqrt(b^2-4ac)/2a x=-(-16)+-sqrt(-16^2-4(4)(-1)/2(4) x=16+-sqrt(256-4(-4)/8 x=16+-sqrt(256+16)/8 x=16+-sqrt(272)/8 and now I'm stuck... help?

OpenStudy (anonymous):

I think the next step would be: x=16+-sqrt(17)*sqrt(16)/8 then x=16+-4sqrt(17)/8 ??

OpenStudy (anonymous):

\[-16t^{2} + 64t + 4 = 0\] would give us \[t = 2 - \frac{ \sqrt{17} }{ 2 }\] and \[t = \frac{ 1 }{ 2 }(4+\sqrt{17})\]

OpenStudy (anonymous):

So its the second value!

OpenStudy (anonymous):

Can you show me how you got that? (: